A label tells you the following: One serving contains 24 grams of total carbohydrate. There are 7 grams of added sugars and 10 grams of total sugars. Each serving supplies two grams of fiber. Estimate the grams of starch in one serving of this food. Show your work and make sure and use proper units of measure
Answer #1

Total carbohydrates in one serving = 24 grams.

Added sugars = 7 grams

Total sugars = 10 grams

Fiber per serving = 2 grams

Grams of starch in one serving of this food = 12 grams

Net carbs in a serving = total carbs - fiber = 24 - 2 = 22 grams

Net carbs = sugars + starch

22 = 10 + starch

Grams of starch in one serving of this food = 22 -10 = 12 grams